ASP:

Ethanol reacts with phosphorus trichloride to yield chloroethane and phosphorous acid according to the following balanced chemical equation: 3 C2H5OH + PCl3 → 3 C2H5Cl + H3PO3 If a chemist starts with 105.0 g of ethanol and 135.65 g of phosphorus trichloride, then what is the limiting reagent?

Vocaloid:

you have 2 reagants, 1 is limiting and the other is not. the limiting reagent will produce fewer moles of product, when compared to the non-limiting reagent. therefore, you can simply set up the stoichiometry to convert g of reagent to moles of product. either product is fine for the calculation, as long as you're consistent. so for these calculations, I will see whether 105.0 g of ethanol (C2H5OH) or 135.65 g of phosphorus trichloride (PCl3) will produce less H3PO3 \[\frac{ 105.0g~ethanol }{ }\frac{ 1~mol~ethanol }{ 46.07g~ethanol }\frac{ 1~mol~H3PO3 }{ 3~mol~ethanol }\] repeat the setup for 135.65 g of phosphorus trichloride and see whether the ethanol or the phosphorus trichloride results in less product